Potassium hydrogen phthalate is a solid, monoprotic acid frequently used in the laboratory to standardize strong base solutions. It has the unwieldy formula of KHC8H4O4. This is often written in shorthand notation as KHP.
If 5.15 grams of KHP are needed to exactly neutralize 32.2 mL of a barium hydroxide solution, what is the concentration of the base solution?

ANSWER: ______ M barium hydroxide
Answer #1

To solve this question, first we need to write balance chemical reaction between KHP and Ba(OH)2 which is given by

2 KHC8H4O4 + Ba(OH)2 ---> Ba(KC8H4O4)2 + 2 H2O

Now find out the moles of KHP first.

5.15 g KHC8H4O4 * (1 mol KHC8H4O4/204.22 g KHC8H4O4) = 0.025217 mol KHC8H4O4

Now we need to find moles of Ba(OH)2. For this check mole to mole ratio for KHC8H4O4 to Ba(OH)2 which is 2:1.

2 mol KHC8H4O4 = 1 mol Ba(OH)2

0.025217 mol KHC8H4O4 = 0.025217 mol KHC8H4O4 * (1 mol Ba(OH)2/2 mol KHC8H4O4) = 0.0126085 mol Ba(OH)2

Now we have moles of Ba(OH)2 and its volume. To find molarity we use the formula

Molarity = moles/Litre of solution

First convert volume of Ba(OH)2 in litre

32.2 ml * (1L/1000 ml) = 0.0322 L

Molarity = 0.0126085 mol / 0.0322 L = 0.3915 mol/L or M.

Therefore molarity of Ba(OH)2 solution will be 0.392 M.

Thanks.